gu.lang <@t> gmx.at
Thu Feb 19 14:15:38 CST 2009
Hi dear listmembers!
What troubleshooting would you suggest for this problem:
Her2neu FISH with Vysis-standard protocol (30 min 80°C pretreatment
solution, 10 min 37°C 0,05% pepsin)
The staining result is clear CEP-staining in green but no LSI-staining in
With the same run, I had an amplified case, that was clearly seen with more
or less bright signals.
More information about the Histonet